# Galley: Modern Query Optimization for Sparse Tensor Programs

### v3

@@ 7.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g. non-zero) entries. The uniform estimator uses only this statistic and relies on the assumption that these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the size of the index attribute’s active domain is equal to the size of the dimension (Selinger et al., 1979).
@@
Given a tensor $A_{i_{1},\ldots,i_{k}}\in\mathbf{R}^{n_{i_{1}}\times\ldots\times n_{i_{k}}}$, we simply count the non-fill values in the tensor, $nnz(A)$, and note the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
@@
To handle an annihilating pointwise operation, we calculate the probability that a point in the output was non-fill in all inputs, then multiply this with the dimension space of the output. For a set of inputs $A^{(1)}_{I_{1}}\ldots A^{(l)}_{I_{l}}$ and output $C_{I_{C}}$ where each $I_{j}$ is a set of indices, this probability is
@@
To handle an non-annihilating pointwise operation, we calculate the probability that an entry in the output was fill in all inputs. Then, we take the compliment to get the probability that it was non-fill in all inputs and multiply this with the output dimension space. Using the notation from above,
@@
Given an input tensor $A_{I}$ which we are aggregating over the indices $I^{\prime}$, we compute the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor wasn’t fill.
$nnz(C)\approx\left(\prod_{i\in I\setminus I^{\prime}}n_{i}\right)\cdot\left(1-\left(1-\frac{nnz(A_{I})}{\prod_{i\in I}n_{i}}\right)^{\prod_{i\in I^{\prime}}n_{i}}\right)$
@@
The estimation function simply returns the cardinality statistic stored about the current tensor.
7.3.2. Degree Statistics & The Chain Bound
We keep degree statistics as the default in Galley, and we use them to compute upper bounds on the number of non-fill entries in intermediate expressions. A degree statistic, denoted $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, if you have a matrix $A_{ij}$, then $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding which has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
@@ 7.3.2. Degree Statistics & The Chain Bound
We first compute the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, we sum over the $X$ dimensions and take the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$
@@
Annihilating map operations function as conjunctive queries with respect to the sparsity patterns of the inputs. Therefore, any degree statistics that are valid for an input are also valid about the output, so we compute the statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by a union,
@@
In this case, we need to be more careful to ensure that we maintain our upper bounds. First, we extend the degree constraints from each input to cover the full set of indices. For example, if we have $DC_{X}(i|j)$ and want to extend it to the dimension $k$, then we compute $D_{X}(ik|j)=DC(i|j)*n_{k}$. Then, we compute degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition,
@@
We calculate our upper bound (eq. perform sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Our search begins with the empty set, then we use breadth first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bound the number of non-zeros in the result.
8. Additional Optimizations
In this section, we outline a few additional techniques we apply to improve the optimization and execution of tensor programs.
8.1. Just-In-Time Physical Optimization
When optimizing a large program, estimates of sparsity become less reliable because they require more steps of inference on top of grounded inputs. This poses a challenge for the physical optimizer as it estimates the cost of different loop orders, formats, and access protocols. To ameliorate this, Galley performs a very simple form of adaptive optimization by waiting to perform physical optimization on each logical query until all of its aliases have already been executed. This allows it to calculate the size of prior intermediate results and update the relevant statistics before performing physical optimization. This makes that cost estimates in that optimization more accurate and produces a better physical plan.
8.2. Common Sub-Expression Elimination
We take a straightforward approach to avoiding redundant computation. Before executing each physical query, we first canonicalize the right hand side, hash it, and compute its hash value. We use this to check a cache to see if it has been executed already. If so, we immediately return the value from the cache rather than re-executing the query. Similarly, after computing a query, we store the result in this cache for later reuse.
Figure 6. Machine Learning Algorithms Over Joins
@@ 9. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) machine learning algorithms, (2) subgraph counting, and (3) breadth-first search. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while only requiring a concise, declarative input program from the user. Overall, we show that Galley:
@@
  • Selects optimal tensor formats for intermediates, beating both the dense and sparse formats for 3/5 graphs
Experimental Setup
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://github.com/kylebd99/Galley. We used the sparse tensor compiler Finch for execution, and all methods are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 7.3.2 for sparsity estimation. Experiments for all methods are run three times and the minimum execution time is reported. This removes the compilation overhead which we address separately in Fig. 9.
@@ 9.1. Machine Learning Algorithms
To explore end-to-end program optimization, we experiment with simple machine learning algorithms over joins, represented entirely in tensor algebra. For this, we use the TPC-H benchmark and consider two join queries, a star query and a self-join query. First, we perform a star join over the line items table to bring together features about suppliers, parts, orders, and customers. This is expressed as follows where $L,S,P,O,$ and $C$ are tensors representing the lineitems, suppliers, parts, orders, and customers tables, respectively.
@@
The non-zero values in $S,P,O$ and $C$ are disjoint along the $j$ axis, so the addition in this expression serves to concatenate features from each source, resulting in 139 features after one-hot encoding categorical features. The self-join query compares line items for the same part based on part and supplier features. In this case, the feature data is a 3d tensor because the data points are keyed by pairs of line items.
@@
Given these definitions, we consider a range of ML algorithms; 1) linear regression inference 2) logistic regression inference 3) covariance matrix calculation and 4) neural network inference. As a comparison point, we also implement two versions of each of these using the Finch compiler. The dense version uses a fully dense matrix to represent the feature matrix whereas the sparse version uses a sparse level for the features to compress the one-hot encoded features. Fig. 6 shows that Galley is 2100×2-100$\times$2 - 100 × faster for all but one experiment. This performance improvement is derived from Galley pushing the computation into the definition of $X$. For example, linear regression inference is fundamentally a matrix-vector multiplication, and Galley pushes the vector multiplication all the way down to the feature matrices $S,P,O,$ and $C$. Note that we do not provide a dense implementation for the self-join query because it runs out of memory due to the highly sparse $i_{1}$, $i_{2}$ dimensions.
@@
Figure 10. BFS Runtime
@@ 9.2. Subgraph Counting
To test Galley’s performance on large, sparse problems, we implement a few common sub-graph counting benchmarks. The conversion from sub-graph counting to sparse tensor algebra is straightforward. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$, then we can represent the count as,
@@
To handle vertex labels, we add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study”(Sun and Luo, 2020; Park et al., 2020). We restrict the latter benchmark to query graphs with up to 8 vertices, hence the ”_lite” suffix . Because this is a relational workload, we compare with DuckDB, a modern OLAP database (Raasveldt and Mühleisen, 2019). To tease apart the impact of logical optimization from physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. Lastly, we provide results for the greedy logical optimizer as well.
In Fig. 7, we see median execution times which are up to 100×100$\times$100 × lower when comparing Galley to DuckDB. When using Galley’s physical optimizer with Finch as opposed to using DuckDB as an execution engine, we still see up to 10×10$\times$10 × lower median execution time. The former shows the benefit of Galley’s logical optimizer and the variable elimination framework. The latter demonstrates the benefits of using STCs which enable a wide range of formats and produce highly efficient code. Further, DuckDB hits the 300 second timeout on 196 out of 400 queries in the youtube_lite benchmark. This is reduced to 5 and 18 timeouts when using Galley with a DuckDB and Finch execution engine, respectively. The greedy optimizer performs similarly to the exact optimizer on these workloads with the latter providing improvements on the slowest queries. Fig. 8 shows the mean optimization time for each of the methods on each of the workloads. Galley has a mean optimization time of less than .15.15.15.15 seconds across all workloads, approaching the time taken by the highly efficient DuckDB optimizer.
@@
Lastly, because it performs compilation using Finch at runtime, Galley incurs a compilation latency when it first requests each tensor kernel implementation. Fortunately, these kernels are cached automatically by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 9. On the simpler workloads which often reuse kernels, this cost is minimal. However, the more complex workloads both reuse kernels less and require compiling more complex kernels, significantly increasing compilation time. This suggests a need for faster STC compilation or the use of an interpreted engine for complex queries.
9.3. Breadth First Search
@@
To demonstrate the importance of selecting optimal formats, we implement a simple push-based breadth first search algorithm using Galley and different hand coded Finch implementations. Both systems are provided with a single iteration at a time and the total execution time across all iterations is reported. This means that the core optimization question is how to represent the vector of visited vertices and the vector of frontier vertices. The former grows monotonically over the course of the algorithm while the number of non-zeros in the latter forms a curve, peaking in the middle iterations. We provide two implementations of Finch 1) using a sparse vector for both intermediates and 2) using a dense vector for both intermediates. Fig. 10 shows that Galley is faster than both dense and sparse approaches for 3 of the 5 graphs and is competitive for all graphs. Note that this includes the optimization time for Galley.
10. Conclusion & Limitations

### v4

@@ 7.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g., non-zero) entries. The uniform estimator uses only this statistic and relies on the assumption that these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the size of the index attribute’s active domain equals the size of the dimension (Selinger et al., 1979).
@@
Given a tensor $A_{i_{1},\ldots,i_{k}}\in\mathbf{R}^{n_{i_{1}}\times\ldots\times n_{i_{k}}}$, this function simply counts the non-fill values in the tensor, $nnz(A)$, and notes the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
@@
To handle an annihilating pointwise operation, this function calculates the probability that a point in the output was non-fill in all inputs, then multiplies this by the dimension space of the output. For a set of inputs $A^{(1)}_{I_{1}}\ldots A^{(l)}_{I_{l}}$ and output $C_{I_{C}}$, where each $I_{j}$ is a set of indices, this probability is
@@
To handle an non-annihilating pointwise operation, this function calculates the probability that an entry in the output was fill in all inputs. Then, it takes the compliment to get the probability that it was non-fill in all inputs and multiplies this by the output dimension space. Using the preceding notation:
@@
Given an input tensor $A_{I}$ to aggregate over the indices $I^{\prime}$, this function computes the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor was not fill:
$nnz(C)\approx\left(\prod_{i\in I\setminus I^{\prime}}n_{i}\right)\cdot\left(1-\left(1-\frac{nnz(A_{I})}{\prod_{i\in I}n_{i}}\right)^{\prod_{i\in I^{\prime}}n_{i}}\right).$
@@
The estimation function simply returns the current tensor’s stored cardinality statistic.
7.3.2. Degree Statistics and the Chain Bound
Galley stores degree statistics as the default, and it uses them to compute upper bounds on the number of non-fill entries in intermediate expressions. A degree statistic, denoted as $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, given a matrix $A_{ij}$, then $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding that has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
@@ 7.3.2. Degree Statistics and the Chain Bound
This function first computes the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, it sums over the $X$ dimensions and takes the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$.
@@
Annihilating map operations serve as conjunctive queries with respect to the sparsity patterns of the inputs. Therefore, any degree statistics that are valid for an input are also valid about the output. This function computes the statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by a union
@@
In this case, the function must be especially careful to ensure that it maintains the upper bounds. First, it extends the degree constraints from each input to cover the full set of indices. For example, if we have $DC_{X}(i|j)$ and want to extend it to the dimension $k$, the function computes $D_{X}(ik|j)=DC(i|j)*n_{k}$. Then, it computes degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition:
@@
This function calculates an upper bound (eq. performs sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Its search begins with the empty set; it then uses a breadth-first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bounds the number of non-zeros in the result.
Figure 6. ML Inference Over Joins
@@ 8. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) ML algorithms over joins, (2) subgraph counting, and (3) breadth-first search. We choose those workloads because they exercise different aspects of our optimizer on real-world use-cases: ML algorithms over joins require careful logical optimizations over programs with mixtures of dense and sparse inputs and non-linear operators; subgraph counting requires both logical and physical optimization of complex sum-product expressions over highly sparse inputs and demonstrates Galley’s advantage over a relational engine even for very sparse workloads; breadth-first search requires careful selection of tensor formats over the course of the computation, showing the benefit of physical optimization for even simple computations. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while requiring only a concise, declarative input program from the user. Overall, we show that Galley:
@@
  • Selects optimal formats for intermediates, outperforming both fully dense and sparse formats for 4/5 graphs in a BFS application.
Experiment Setup
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://anonymous.4open.science/r/Galley-21BF/. We used the sparse tensor compiler Finch [^2] https://github.com/FinchTensor/Finch.jl for execution, and all experiments are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 7.3.2 for sparsity estimation. Experiments for all methods are run five times, and the mean execution time is reported. We perform all experiments on a warm cache, and we separately report the compilation and optimization times.
@@ 8.1. Machine Learning Algorithms
To explore end-to-end program optimization, we experiment with simple ML algorithms over joins, represented entirely in tensor algebra. For this, we use the TPC-H benchmark at a scale factor of .25.25.25.25 in order to keep the comparison methods in memory [^3] Galley’s execution tends to be less memory-intensive, and we have successfully run it at scale factor 1.5.. We consider two join queries on this dataset: star and self-join. First, we perform a star join over the line items table to gather features about suppliers, parts, orders, and customers. This is expressed as follows, where $L,S,P,O,$ and $C$ are tensors representing the line items, suppliers, parts, orders, and customers tables, respectively:
@@
The non-zero values in $S,P,O$ and $C$ are disjoint along the $j$ axis, so the addition in this expression serves to concatenate features from each source, resulting in 139 features after one-hot encoding categorical features. The self-join query compares line items for the same part based on part and supplier features. In this case, the feature data is a 3D tensor because the data points are keyed by pairs of line items:
@@
Given these definitions, we consider a range of ML algorithms: (1) linear regression inference, (2) logistic regression inference, (3) covariance matrix calculation, and (4) neural network inference. For comparison, we also implement two versions of each of these using the Finch compiler. The dense version uses a fully dense format to represent the features, whereas the sparse version uses a sparse level for the features to compress the one-hot encoding. We also implemented these algorithms using the Pandas and Numpy libraries for reference.
These algorithms stress the ability of Galley to handle both combinations of complex operators and combinations of sparse and dense inputs. The definitions of the feature tensors combine pointwise multiplication and addition, and algorithms like logistic regression and neural networks wrap these definitions in more non-linear operators (e.g. relu and sigmoid) and aggregates. Further, while the line item tensor is highly sparse, both the feature and parameter tensors are moderately to fully dense. In all of these examples, Galley is able to manipulate these complex expressions to avoid materializing large, sparse intermediates. Instead, its logical optimizer aggressively pushes the computation and aggregation down to the smaller, denser inputs. Consider logistic regression over the star query feature matrix. In this case, Galley distributes the parameter vector over the addition and performs the summations $\sum_{j}S_{sj}\theta_{j},\sum_{j}P_{pj}\theta_{j},\sum_{j}O_{oj}\theta_{j},$ and $\sum_{j}C_{cj}\theta_{j}$ before handling the remaining operation in one kernel. By doing this, it materializes a dense vector at every point and removes the feature dimension before operating on the large, sparse line item tensor.
Execution Time. Fig. 6 shows that the execution time of Galley’s optimized programs is .5300×.5-300$\times$.5 - 300 × faster than the sparse Finch implementation. For the regression/neural network problems, this stems from pushing the parameter vector/matrix down to the feature matrices as discussed above. For the covariance calculation over the self join, Galley fully distributes the multiplication over the addition. Each distributed term is the product of four copies of the lineitem table $l_{i_{1},s_{1},p_{1}},l_{i_{2},s_{2},p_{1}}l_{i_{1},s_{3},p_{2}}l_{i_{2},s_{4},p_{2}}$ and one feature matrix from each $X$, e.g. $S_{s_{1},j}P_{p_{1},k}$. Within each term, Galley performs the summation over $i_{1}$ and $i_{2}$ first, before handling the summation of $p_{1}$ or $p_{2}$. Because $i_{1}$ and $i_{2}$ are keys, unlike $p_{1}$ and $p_{2}$, these initial sums produce small, dense intermediates, speeding up the whole computation. Lastly, the range of Galley’s execution times for the five iterations of each experiment was never more than $25\%$ of the mean.
Optimization Time. Fig.  6 also shows that Galley’s optimizer has a reasonable overhead in this setting. For all experiments except for Covariance (SQ), Galley is faster than competing methods even when including the optimization time. Concretely, optimization takes .$15-3.0$ seconds on these workloads. The covariance calculations take longer to optimize due to the number of terms produced when they are fully distributed.
@@
Figure 10. Sparsity Estimator Comparison
@@ 8.2. Subgraph Counting
In this section, we test Galley’s ability to optimize programs with a large number of highly sparse inputs. To do this, we implement a few common sub-graph counting benchmarks. The conversion from sub-graph counting to sparse tensor algebra is straightforward and results in traditional einsum expressions. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$; we can represent the count as
@@
To handle vertex labels, we add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study”(Sun and Luo, 2020; Park et al., 2020). We restrict the latter benchmark to query graphs with up to 8 vertices (hence the ”_lite” suffix). Because this is a relational workload, we compare it with DuckDB, a modern OLAP database (Raasveldt and Mühleisen, 2019). To separately discern the impact of logical vs physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. Finally, we provide results for the greedy logical optimizer, as well.
Logical Optimization. Fig. 7 shows the execution time for each method and benchmark. The comparison between ‘DuckDB‘ and ‘Galley + DuckDB Backend‘ demonstrates the benefits of Galley’s logical optimizer. Galley’s logical optimizer breaks down the program into a series of aggregations which effectively minimize the necessary computation and materialization, resulting in a dramatic performance improvement. This has the largest impact on graphs with high degree values and/or high skew like the social network graphs, ‘dblp_lite‘ and ‘youtube_lite‘. In these cases, not pushing down aggregates can result in very large intermediate results. On low-degree graphs like ‘aids‘, these optimizations have less impact. Further, DuckDB hits the 300 second timeout on 50 out of 120 queries in the youtube_lite benchmark. In contrast, Galley never times out across all workloads.
Physical Optimization. The impact of Galley’s physical optimizer can be seen by comparing ‘Galley‘ with ‘Galley + DuckDB Backend‘. By producing efficient STC kernels, Galley’s median execution is up to 8x faster than DuckDB even with the same logical plan. This shows that Galley is selecting efficient loop orders and formats which allows it to take advantage of the highly flexible language of STCs.
Optimization Time. Fig. 8 shows the mean optimization time for each method on each workload. Galley has a mean optimization time of less than .15.15.15.15 seconds across all workloads, approaching the time taken by the highly efficient DuckDB optimizer.
@@
Compilation Time. Because it performs compilation using Finch at runtime, Galley incurs a compilation latency the first time that it invokes each unique STC kernel. Fortunately, these kernels are automatically cached by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 9. On the simpler workloads, which often reuse kernels, this cost is minimal. However, the more complex workloads both reuse kernels less and require compiling more complex kernels, significantly increasing compilation time. This suggests a need for faster STC compilation or the use of an interpreted engine for complex queries.
Sparsity Estimation. Finally, in Fig. 10, we use the sub-graph matching workloads to compare sparsity estimators and their effect on performance. Across all workloads, we see that the chain bound significantly out-performs the uniform estimator. This is because graph datasets tend to be skewed and correlated. Vertices have drastically different edge counts, and vertices with many edges tend to connect to vertices with many edges. This violates the implicit assumptions of the uniform estimator, leading it to optimistically materialize large intermediates. Due to this, the uniform estimator runs out of memory on many queries in the ‘youtube_lite‘ workload, so it is not pictured in this figure.
Figure 11. BFS Execution Time
8.3. Breadth-First Search
@@
To demonstrate the importance of selecting optimal physical formats for tensors, we implement a simple breadth-first search algorithm using Galley and different hand-coded Finch implementations. Both systems receive a single iteration at a time, and the total execution time across all iterations is reported. The core computation for each iteration is a masked sparse matrix times sparse vector operation that computes the new frontier vector. The main optimization decision is the visited and frontier vectors’ formats. The former grows monotonically over the course of the algorithm, while the number of non-zeros in the latter forms a curve, peaking in the middle iterations. We provide two implementations of Finch, using either a sparse or a dense vector for both intermediates. Fig. 11 shows that Galley is significantly faster than both dense and sparse approaches for 4 of the 5 graphs and is competitive for all graphs. During a single run of BFS, Galley begins with a sparse visited vector before switching to bytemap and dense formats as more nodes are discovered. The frontier vector begins as sparse, becomes dense, then again becomes sparse for the last few iterations as most nodes are deactivated. For 4/5 graphs, the total optimization time for Galley (not depicted in the figure) is less than .25.25.25.25 seconds. RoadNet has a very wide diameter leading to many iterations and a total optimization time of $4$ seconds. This experiment demonstrates the utility of sparsity-aware format selection, and future work should consider ways to amortize optimization time for iterative workloads.
9. Conclusion and Limitations